Determine k so that the line through (k+2,k) and (3,2) has a slope of 3.
Answer by nerdybill(7384)

Determine k so that the line through (k+2,k) and (3,2) has a slope of 3.
.
Since slope is defined as:
(y2 - y1)/(x2 - x1)
we have
3 = (2-k)/(3-(k+2))
3 = (2-k)/(3-k-2)
3 = (2-k)/(1-k)
3(1-k) = (2-k)
3-3k = 2-k
3 = 2+2k
1 = 2k
1/2 = k